A saturated solution of PbI2 has a concentration of 1.65 × 10-3 mol/dm3
.
i. Write the reversible reaction for PbI2 breaking into its constituent ions.
ii. Write the equilibrium expression (Ksp).
iii. Calculate the concentration of each ion, Pb2+ and I-
.
iv. Calculate the Ksp.
1
Expert's answer
2020-12-08T05:29:58-0500

(i) PbI2 = Pb2+ + 2I-

(ii) Ksp = [Pb2+]*[I-]2

(iii) [Pb2+] = c(PbI2) = 1.65*10-3 mol/L

[I-] = 2*c(PbI2) = 2*1.65*10-3 mol/L = 3.3*10-3 mol/L

(iv) Ksp = [Pb2+]*[I-]2 = (1.65*10-3)*(3.3*10-3)2 = 1.7968*10-8